OpenStudy (anonymous):

15t/5 = 2t +3/6 solve the proportion

OpenStudy (anonymous):

Please use brackets. I think you mean 15t/5 = (2t+3)/6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First step is pretty obvious 15t/5 = (2t+3)/6 3t = (2t+3)/6 Now multiply both sides by 6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then